# A Review of Design Concerns in Superconducting Quantum Circuits

### v4

@@ I Introduction
    Fabricate the device, measure its behavior, and compare to the simulation output (which matches target behavior). If experiment results deviate significantly from simulation, refine the simulation and/or fabrication pipelines and repeat steps 3-6.
The review is structured according to this design loop workflow. It begins with a brief description of the physics of superconducting quantum circuits at an effective Hamiltonian level. We then discuss depictions of these circuits as being composed of lumped elements in a particular circuit graph, and how these depictions can be used to design the desired effective Hamiltonian. We follow this by discussing how a physical device layout is designed and how the layout can effect behavior. We then discuss simulations that can be used to predict device behavior given a layout. We describe how the device can change between layout and fabrication. And finally we close the loop by describing how measured device behavior can be used to improve the next round of designs. Throughout each discussion we describe how the steps interact with each other. The goal of this review is not to provide a comprehensive, detailed description of each aspect of design—in many cases these articles have already been written, and we refer the reader to them throughout the text. Instead this review is intended as an overview that can be used as an introduction to superconducting device design for new researchers in the field, and a centralized reference that they can use to find more detailed resources on each aspect of the problem. LABEL:tab:references summarizes these references, organized by topic.
@@ II Superconducting Qubits
Here we give a brief introduction to the behavior of superconducting circuits. More comprehensive reviews can be found in references [2, 3, 4, 5, 1, 6, 7, 8, 9].
Figure 1: Diagram of an inductor (left, orange), Josephson junction (center, blue), and capacitor (teal, right), forming a parallel LJC circuit. Each element has an associated energy. Tiling LJC subcircuits in series and parallel allows one to build up an arbitrary circuit.
A circuit can be modeled as a set of nodes (each with its own voltage) connected by branches (each carrying its own current). A branch can be formed by any lumped circuit element. Neglecting loss and quantum phase slips (see Section III.1), in a superconducting circuit such an element can be an inductor (with inductance $L$), a Josephson junction (with critical current $I_{0}$), or a capacitor (with capacitance $C$) (we neglect distributed elements for the moment). One can combine these into the parallel LJC subcircuit. See Fig. 1 for a diagram. Note that the capacitance here may be entirely due to the Josephson junction geometry or may have some additional contribution from a physical shunting capacitor, and the inductance is a physical shunting inductance in parallel with the junction. This is the simplest possible circuit that captures all possible branches between two nodes. It produces the Hamiltonian
@@
Here $\hat{n}$ is the charge number operator (where charge is quantized in units of Cooper pairs, i.e., $2e$), $n_{g}$ is a constant that accounts for the effect of voltage bias across the capacitor, $\hat{\phi}$ is the dimensionless gauge-invariant node flux (that is, the phase difference across the junction/inductor–see Section III.1 for a rigorous definition), and $\phi_{ext}$ is the dimensionless external flux threading the junction-inductor loop $\phi_{ext}=\frac{\Phi_{ext}}{\varphi_{0}}$, where $\varphi_{0}\equiv\hbar/2e$ is the reduced magnetic flux quantum. Note that node flux is distinct from flux threading a loop, as explained in Section III.1. The behavior of this subcircuit can be entirely captured by the component energies $E_{c}\equiv e^{2}/2C$, $E_{J}\equiv\varphi_{0}I_{0}$, and $E_{L}\equiv\varphi_{0}^{2}/L$, along with the environmental bias variables $n_{g}$ and $\phi_{ext}$. A time-dependent $\phi_{ext}$ can lead to another term, but we will ignore this for simplicity [10, 11, 12].
@@
While the parallel LJC subcircuit need not be a qubit, one can produce certain qubit types from a single LJC circuit by varying the energies. This includes the ubiquitous transmon ($E_{L}=0$, $E_{J}\gg E_{c}$) and the many varieties of fluxonium ($E_{c}\sim E_{J}\gg E_{L}$), as well as harmonic oscillators ($E_{J}=0$). Subcircuits such as this can be placed in series or parallel to generate an arbitrary lumped element circuit—omission of an element can be represented as setting its energy to 0. Recent work has enumerated all possible combinations of such 2-node building blocks into circuits with up to 5 nodes [13], providing pre-solved Hamiltonians for such circuits. General circuit Hamiltonians can be built up as described in Section III.1. We have omitted loss (i.e., resistance) from this picture, which we will discuss later in Section IV.4.
@@
This circuit Hamiltonian description is accurate and, as we will see, is the level at which device designers must work. Unfortunately it is often inconvenient to work with when designing quantum architectures, algorithms, and gates, and for most quantum information science theory. Instead, the circuit Hamiltonian is typically reduced to an effective Hamiltonian description, treating the circuit as a set of modes of oscillation which are best described as qubits, qudits, harmonic oscillators, and weakly anharmonic oscillators. These modes are coupled, typically with couplings that are weak enough that the original modes still form a useful basis. As an illustrative example, consider a single transmon coupled to a linear $LC$ resonator via a coupling capacitor as in Fig. 2. The circuit-level Hamiltonian is
@@
where the $T$ and $R$ subscripts refer to transmon and resonator circuit elements and operators, respectively, and the charge number operators $n_{T}$, $n_{R}$ are the charge differences from the highlighted nodes to ground. This can be represented by an effective Hamiltonian:
@@
This depiction is conceptually simple, but is often not accurate enough for, e.g., designing gate schemes [16, 17, 18], understanding measurement-induced state transitions [19, 20], or interpreting spectroscopic results [21, 22] (see below). One message we hope the reader takes away from this review is: do not neglect the higher transmon levels or the non-RWA terms unless you have already confirmed they will not affect the result! This means that one must first calculate behavior including higher levels and non-RWA terms and compare to the behavior without. Then, if there is no significant difference, one could neglect the higher levels in any situation where the qubit approximation / RWA is expected to work even better. For higher levels, this means only using the qubit approximation in cases where the ratio of the relevant energy to the anharmonicity (e.g., the ratio of Rabi frequency to anharmonicity in a single-qubit gate, or the ratio of coupling strength to anharmonicity in a two-qubit coupling Hamiltonian) is even smaller than in the case where the approximation has been validated. For the RWA, this means only using it in cases where the ratio of difference between the relevant energies to their sum (e.g. the qubit and resonator if calculating energy shifts, or the drive frequency and qubit if calculating gate behavior) is even smaller than in the case where the approximation has been validated.
@@
where $\Sigma\equiv\omega_{R}+\omega_{T}$. Note that the $\Sigma$ terms arise due to the non-excitation-conserving couplings $ab$ and $a^{\dagger}b^{\dagger}$, and are omitted in many analyses. However, they often can be significant, especially when computing the Lamb shift $\chi_{L}\equiv\omega_{R}^{\prime}-\omega_{R}$. A typical parameter regime is $\omega_{R}/2\pi\approx 7$ GHz, $\omega_{T}/2\pi\approx 4$ GHz, and so ommitting the $\Sigma$ term overestimates $\chi_{L}$ by approximately 25%. Again, these expressions were derived perturbatively, and a fuller analytical or numerical treatment is required to capture features such as the small inherited resonator self-Kerr term $\eta_{R}$ [23], the precise value of the transmon self-Kerr $\eta_{T}$ [14], and higher-order nonlinearities. These features are quite crucial when fine-tuning gate parameters or using the transmon to control a cavity qubit encoding [24, 25, 26, 27]. However, the goal of this section is merely to give an overview of the physics of these circuits, so we neglect these higher-order effects.
@@ III Circuit Graph and Parameters
It is possible to perform an electromagnetic simulation of a device layout and directly extract its energy spectrum and eigenstates, leading directly to an effective Hamiltonian description [28] (see Section V.5 for an in-depth discussion). However, if one wishes to analytically derive the eigenstates and energies, develop a fundamentally new type of qubit, or sweep over a variety of device designs, such simulations are not appropriate. Instead, researchers describe the device by its circuit graph, treating it as a set of lumped capacitors, inductors, and junctions between circuit nodes. A general example is shown in Fig. 3. This circuit graph description provides a useful middle step between the physical layout and the diagonalized Hamiltonian. This is an artificial creation, as only the device layout and the behavior it implements are actually physical. However, the circuit graph is such a useful tool that we will treat it as if it were physical.
Figure 3: An arbitrary circuit graph. Nodes (black dots) are connected by branches which may be inductive, capacitive, or Josephson elements. Each loop has an independent bias flux; each capacitive branch also has a bias voltage, which we omit for simplicity. From the set of capacitive branches one can define a spanning tree and thus define the node fluxes and charges.
@@ III.1 From Graph to Hamiltonian
First, we must write a circuit graph so that any connected nodes always have a capacitive branch between them, as in Fig. 3. This is physically accurate, as any parts of a physical device that are close enough to couple inductively or through a junction will also have some stray capacitance. From this graph of $N$ nodes we define the $N\times N$ capacitance matrix $\mathbf{C}$ with off-diagonal elements $-C_{mn}$, the negative of the capacitance between nodes $m$ and $n$. The on-diagonal elements are defined so that each row and column sums to 0, and the matrix is symmetric by default.
@@
We then define a spanning tree of capacitances, which connects every node to ground through a path of only capacitive couplings without forming any loops. This means that every node has a single unique path to ground through a series of capacitive branches. Summing the branch fluxes from ground to the $i$th node thus defines the node fluxes $\Phi_{i}$, which we divide by the reduced flux quantum $\varphi_{0}\equiv\hbar/2e$ to get the dimensionless node flux. Any external bias flux threading a loop is taken to add on branches which are not part of the tree. We can then use node fluxes as our coordinates and write the potential energy in terms of these coordinates. For instance, an inductor $L_{ij}$ connecting nodes $i$ and $j$ will contribute energy $E=(\phi_{i}-\phi_{j}+\phi_{ext,ij})^{2}/2L_{ij}$, where $\phi_{ext,ij}$ is the dimensionless bias flux assigned to this branch of the loop. Likewise, the energy contribution from a Josephson junction depends on the difference in node fluxes across it (plus any bias flux assigned to that part of a loop).
@@
Since we are using node fluxes as our coordinates, we can write the “kinetic” energy in terms of the node fluxes’ time derivatives. The energy contribution from a capacitor connecting nodes $i$ and $j$ is then $E=C(\dot{\phi_{i}}-\dot{\phi_{j}})^{2}/2$, i.e., it depends on the difference of these node flux time derivatives across it. Combining the kinetic and potential energies we can write the Lagrangian $\mathcal{L}$. From this we can define the node charges as the canonical momenta, $q_{n}\equiv\partial\mathcal{L}/\partial\dot{\phi_{n}}$, and write the classical Hamiltonian in terms of node charges and fluxes, plus bias voltages and fluxes. At this point we change charge/flux coordinates to operators and obtain the quantum Hamiltonian as in Eq. 2.
@@
This process sounds daunting even when summarized. Luckily, the procedure is quite straightforward, if complicated, and has been both written about in great detail [29, 30] and has been integrated into open-source codes like scqubits, SQCircuit, and CircuitQ [32, 33, 34]. A user can input an arbitrary circuit, follow the procedure, and obtain a numerical solution of the eigenstates and energies. In practice, this requires truncating the charge and flux basis states so that the Hilbert space is finite and can be diagonalized numerically. However, for a circuit with any significant complexity, a complete numerical solution can be computationally infeasible even when bases are truncated. One solution is to perform hierarchical diagonalization [30, 35, 33]. In this approach, a designer uses their knowledge of a circuit to divide it into subcircuits, each of which has well-defined eigenstates that can be easily solved. The subcircuits are described using a basis of these eigenstates, which again is truncated at some finite dimension. The subcircuits are then coupled together and solved using the product space of these subcircuit basis states. This is similar to a perturbative approach but has the advantage of easily handling degenerate states.
@@
In addition to the challenge of solving an arbitrary circuit, it is typically desirable to have an effective quantized Hamiltonian written in terms of oscillator creation/annihilation operators and Pauli matrices, as in Eq. 3. Such an effective Hamiltonian is more amenable to analytical theory and is usually more useful for quantum gate design and error correction architecture design. One solution is to use knowledge of the intended design to build up subcircuits that are weakly coupled—for instance, the coupled transmon and oscillator in Fig. 2 can be solved as an arbitrary circuit, but it is easier to recognize the transmon and oscillator modes and then add in the coupling in terms of these mode operators. Another approach is to exactly diagonalize the Hamiltonian, then write it in its eigenbasis with a set of Kerr, cross-Kerr, and higher-order terms as in Section II. This approach has the advantage of directly giving the energy spectrum, but is less useful if environmental conditions (flux or charge biases or drive tones) will be changed, as these can completely change the mode structure. Methods for rapidly calculating terms of interest such as inter-qubit couplings and dispersive couplings have been developed [36, 37], making the procedure less computational intensive.
@@
It is possible to connect nonlinear circuit elements in such a way that a Lagrangian cannot be defined via this procedure. For instance, a Josephson junction in parallel with a quantum phase slip junction has a Hamiltonian description but no Lagrangian description. A phase slip junction is the dual of a Josephson junction, with Hamiltonian contribution $H_{Q}=-E_{Q}\cos\left(2\pi\frac{q}{2e}\right)$—the appearance of the charge in the nonlinear term breaks the standard procedure [38]. Likewise, the Lagrangian for a non-reciprocal device cannot be straightforwardly written down since the branch flux would depend on the direction we choose to calculate it [39]. Fortunately, frameworks for dealing with these situations have been formulated in geometric [40, 38, 41] and exactly quantized [42] terms as well as including open-systems effects [39]. We refer the reader to Ref. [41] for an excellent summary of the literature.
III.2 From Layout to Graph

### v5

@@ I Introduction
    Fabricate the device, measure its behavior, and compare to the simulation output (which matches target behavior). If experiment results deviate significantly from simulation, refine the simulation and/or fabrication pipelines and repeat steps 3–6.
The review is structured according to this design loop workflow. It begins with a brief description of the physics of superconducting quantum circuits at an effective Hamiltonian level. We then discuss depictions of these circuits as being composed of lumped elements in a particular circuit graph, and how these depictions can be used to design the desired effective Hamiltonian. We follow this by discussing how a physical device layout is designed and how the layout can effect behavior. We then discuss simulations that can be used to predict device behavior given a layout. We describe how the device can change between layout and fabrication. And finally we close the loop by describing how measured device behavior can be used to improve the next round of designs. Throughout each discussion we describe how the steps interact with each other. The goal of this review is not to provide a comprehensive, detailed description of each aspect of design—in many cases these articles have already been written, and we refer the reader to them throughout the text. Instead this review is intended as an overview that can be used as an introduction to superconducting device design for new researchers in the field, and a centralized reference for locating more detailed resources on each aspect of the problem. LABEL:tab:references summarizes these references, organized by topic.
@@ II Superconducting Qubits
Here we give a brief introduction to the behavior of superconducting circuits. For more detailed discussions, see references [2, 3, 4, 5, 1, 6, 7, 8, 9].
Figure 1: (a) Diagram of an inductor (left, orange), Josephson junction (center, blue), and capacitor (teal, right), forming a parallel LJC circuit. An external flux $\Phi_{\mathrm{ext}}$ threads the loop formed by the L and J. Each element has an associated energy. Tiling LJC subcircuits in series and parallel allows one to build up an arbitrary circuit, including the fixed-frequency transmon (b), harmonic oscillator (c), tunable transmon (d), 3-junction flux qubit (e), and inductively-shunted transmon (f). In examples (b–f) we gray out any circuit element that is omitted, effectively setting its energy to 0, and draw dashed red lines around the LJC subcircuits that combine to make up the overall circuit.
II.1 Quantizing a circuit
A circuit can be modeled as a set of nodes (each with its own voltage) connected by branches (each carrying its own current). A branch can be formed by any lumped circuit element. Neglecting loss and quantum phase slips (see Section III.1), in a superconducting circuit such an element can be an inductor (with inductance $L$), a Josephson junction (with critical current $I_{0}$), or a capacitor (with capacitance $C$) (we neglect distributed elements for the moment). One can combine these into the parallel LJC subcircuit. See Fig. 1 for a diagram. Note that the capacitance here may be entirely due to the Josephson junction geometry or may have some additional contribution from a physical shunting capacitor, and the inductance is a physical shunting inductance in parallel with the junction. This is the simplest possible circuit that captures all possible branches between two nodes and yields the following Hamiltonian
@@ II.1 Quantizing a circuit
Here $\hat{n}$ is the charge number operator (where charge is quantized in units of Cooper pairs, i.e., $2e$), $n_{g}$ is a constant that accounts for the effect of voltage bias across the capacitor, $\hat{\phi}$ is the dimensionless node flux (that is, the phase difference across the junction/inductor with one side defined as 0—see Section III.1 for a rigorous definition), and $\phi_{ext}$ is the dimensionless external flux threading the junction-inductor loop $\phi_{ext}=\frac{\Phi_{ext}}{\varphi_{0}}$, where $\varphi_{0}\equiv\hbar/2e$ is the reduced magnetic flux quantum. Note that the node flux differs from the loop flux threading the circuit loop, as explained in Section III.1. The charge number operator and node flux obey the canonical conjugation relation $\left[\hat{\phi},\hat{n}\right]=i$ similar to position and momentum. The behavior of this subcircuit can be entirely captured by the component energies $E_{c}\equiv e^{2}/2C$, $E_{J}\equiv\varphi_{0}I_{0}$, and $E_{L}\equiv\varphi_{0}^{2}/L$, along with the environmental bias variables $n_{g}$ and $\phi_{ext}$. A time-dependent $\phi_{ext}$ can lead to another term, but we will ignore this for simplicity [10, 11, 12].
@@
While the parallel LJC subcircuit need not be a qubit, one can produce certain qubit types from a single LJC circuit by varying the energies. This includes the ubiquitous transmon ($E_{L}=0$, $E_{J}\gg E_{c}$) and the many varieties of fluxonium ($E_{c}\sim E_{J}\gg E_{L}$), as well as harmonic oscillators ($E_{J}=0$). Subcircuits such as this can be placed in series or parallel to generate an arbitrary lumped element circuit—omission of an element can be represented as setting its energy to 0. Examples of common simple combinations are shown in Fig. 1(b), with omitted elements grayed out but still shown to emphasize this point (note that in all other figures we do not draw elements which are absent from the circuit, as is typical). Recent work has enumerated all possible combinations of such 2-node building blocks into circuits with up to 5 nodes [13], providing pre-solved Hamiltonians for such circuits. General circuit Hamiltonians can be built up as described in Section III.1. We have omitted loss (i.e., resistance) from this picture, which we will discuss later in Section IV.4.
II.2 Quantum vs. classical descriptions
This description assumes the Josephson junction itself is an ideal Josephson element with no resistive transport and only a single cosinusoidal Hamiltonian contribution $\hat{H}_{J}=-E_{J}\cos\hat{\phi}$. This is equivalent to the canonical classical model of a Josephson junction, the resistively capacitively shunted junction (RCSJ) model [14], with two modifications. The first is to take the limit $R\rightarrow\infty$. This limit is appropriate as superconducting quantum circuits are operated deep in the superconducting limit where normal-state transport is negligible. The second modification is more fundamental: the RCSJ model treats the junction as an AC voltage source and treats voltage and current (i.e., charge and flux) as classical coordinates. A full quantum treatment of the junction treats charge and flux as operators to correctly account for the fact that they do not commute, and uses the quantum Hamiltonian to compute state evolution.
Some numerical and theoretical work has demonstrated that the clssical RCSJ model can predict some behavior that had previously been described as purely quantum [15], even including early entanglement demonstrations [16]. It is also very useful for designing classical circuits such as amplifiers and as a starting point for quantum analysis to model the behavior of a composite circuit that functions as a single junction [17]. However, it cannot explain true quantum behavior, which at this point is well-established via experiments such as remote measurement-induced entanglement via one-way interactions [18], Bell violations [19], microwave-optical photon conversion [20], and the many demonstrations of quantum algorithms on superconducting processors. We therefore use a purely quantum treatment of the circuit, which has been fantastically successful at predicting observed behavior.
On the other hand, the assumption of a cosinusoidal Hamiltonian may not be valid, as recent evidence has shown that junctions may not behave as ideal tunnel junctions [21]. Other junctions such as those using semiconducting nanowires [22, 23], 2D seminconductors [24, 25], or other materials may also be used, which have intrinsically non-sinusoidal current-phase relations. One phenomenological approach is to approximate the junction behavior using a finite Fourier series. That is, one uses $N$ Hamiltonian terms of Josephson harmonics, replacing $-E_{J}\cos\hat{\phi}\rightarrow\sum_{m=1}^{N}-E_{J,m}\cos m\hat{\phi}$. This allows the Hamiltonian to be expressed using a finite number of well-behaved analytic terms.
II.3 Effective Hamiltonian description
@@
This circuit Hamiltonian description is accurate and, as we will see, is the level at which device designers must work. Unfortunately it is often inconvenient to work with when designing quantum architectures, algorithms, and gates, and for most quantum information science theory. Instead, the circuit Hamiltonian is typically reduced to an effective Hamiltonian description, treating the circuit as a set of modes of oscillation which are best described as qubits, qudits, harmonic oscillators, and weakly anharmonic oscillators. These modes are coupled, typically with sufficiently weak couplings such that the original modes still form a useful basis. As an illustrative example, consider a single transmon coupled to a linear $LC$ resonator via a coupling capacitor as in Fig. 2. Here and in all following diagrams we only draw circuit elements which are actually present, in contrast with the grayed out elements in Fig. 1. Defining $E_{C,i}=e^{2}/2C_{i}$ and $E_{L}=\varphi^{2}/L_{R}$, the circuit-level Hamiltonian is
@@
where the $T$ and $R$ subscripts refer to transmon and resonator circuit elements and operators, respectively, and the charge number operators $\hat{n}_{T}$, $\hat{n}_{R}$ are the charge differences from the highlighted nodes to ground. In the language of LJC subcircuits this is the Hamiltonian of 3 subcircuits combined: a transmon (with zero inductive energy), a harmonic oscillator (with zero Josephson energy), and a coupling capacitor (with zero inductive or Josephson energy).
This circuit can be described by an effective Hamiltonian:
@@
This depiction is conceptually simple, but is often not accurate enough in practice for, e.g., designing gate schemes [28, 29, 30], understanding measurement-induced state transitions [31, 32], or interpreting spectroscopic results [33, 34] (see below). One message we hope the reader takes away from this review is: do not neglect the higher transmon levels or the non-RWA terms unless you have already confirmed they will not affect the result! This means that one must first calculate behavior including higher levels and non-RWA terms and compare to the behavior without. Then, if there is no significant difference, one could neglect the higher levels in any situation where the qubit approximation / RWA is expected to work even better. For higher levels, this means only using the qubit approximation in cases where the ratio of the relevant energy to the anharmonicity (e.g., the ratio of Rabi frequency to anharmonicity in a single-qubit gate, or the ratio of coupling strength to anharmonicity in a two-qubit coupling Hamiltonian) is even smaller than in the case where the approximation has been validated. For the RWA, this means only using it in cases where the ratio of difference between the relevant energies to their sum (e.g. the qubit and resonator if calculating energy shifts, or the drive frequency and qubit if calculating gate behavior) is even smaller than in the case where the approximation has been validated.
@@
where $\Sigma\equiv\omega_{R}+\omega_{T}$. Note that the $\Sigma$ terms arise from counter-rotating (non-excitation-conserving) couplings such as $ab$ and $a^{\dagger}b^{\dagger}$, which are neglected under the RWA. However, they often can be significant, especially when computing the Lamb shift $\chi_{L}\equiv\omega_{R}^{\prime}-\omega_{R}$. A typical parameter regime is $\omega_{R}/2\pi\approx 7$ GHz, $\omega_{T}/2\pi\approx 4$ GHz, and so omitting the $\Sigma$ term overestimates $\chi_{L}$ by approximately 25%. Again, these expressions were derived perturbatively, and a fuller analytical or numerical treatment is required to capture features such as the small inherited resonator self-Kerr term $\eta_{R}$ [35], the precise value of the transmon self-Kerr $\eta_{T}$ [26], and higher-order nonlinearities. These features are quite crucial when fine-tuning gate parameters or using the transmon to control a cavity qubit encoding [36, 37, 38, 39]. However, the goal of this section is merely to give an overview of the physics of these circuits, so we neglect these higher-order effects.
@@ III Circuit Graph and Parameters
It is possible to perform an electromagnetic simulation of a device layout and directly extract its energy spectrum and eigenstates, leading directly to an effective Hamiltonian description [40] (see Section V.5 for an in-depth discussion). However, if one wishes to analytically derive the eigenstates and energies, develop a fundamentally new type of qubit, or sweep over a variety of device designs, such simulations may not be appropriate. Instead, researchers describe the device by its circuit graph, treating it as a set of lumped capacitors, inductors, and junctions between circuit nodes. A general example is shown in Fig. 3. This circuit graph description provides a useful middle step between the physical layout and the diagonalized Hamiltonian. This is an artificial creation, as only the device layout and the behavior it implements are actually physical. However, the circuit graph is such a useful tool that we will treat it as if it were physical.
Figure 3: An arbitrary circuit graph. Nodes (black dots) are connected by branches which may be inductive, capacitive, or Josephson elements. Each loop has an independent bias flux; each capacitive branch also has a bias voltage, which we omit for simplicity. From the set of capacitive branches one can define a spanning tree - effectively choosing a gauge for the flux variables - and thus assign node fluxes and charges accordingly.
@@ III.1 From Graph to Hamiltonian
First, we must write a circuit graph so that any connected nodes always have a capacitive branch between them, as in Fig. 3. This is physically accurate, as any parts of a physical device that are close enough to couple inductively or through a junction will also have some stray capacitance. From this graph of $N$ nodes we define the $N\times N$ capacitance matrix $\mathbf{C}$ with off-diagonal elements $-C_{jk}$, the negative of the capacitance between nodes $j$ and $k$. The on-diagonal elements are defined so that each row and column sums to 0, and the matrix is symmetric by default.
@@
We then define a spanning tree of capacitances, which connects every node to ground through a path of only capacitive couplings without forming any loops. This means that every node has a single unique path to ground through a series of capacitive branches. Summing the branch fluxes from ground to the $j$th node thus defines the node fluxes $\Phi_{j}$. Any external bias flux threading a loop is taken to add on branches which are not part of the tree. We can then use node fluxes as our coordinates and write the “potential” energy in terms of these coordinates. For instance, an inductor $L_{jk}$ connecting nodes $j$ and $k$ will contribute energy $E=(\Phi_{j}-\Phi_{k}+\Phi_{ext,jk})^{2}/2L_{jk}$, where $\Phi_{ext,jk}$ is the bias flux assigned to this branch of the loop. Likewise, the energy contribution from a Josephson junction depends on the difference in node fluxes across it (plus any bias flux assigned to that part of a loop), $E=-E_{J}\cos\left[(\Phi_{j}-\Phi_{k}+\Phi_{ext,jk})/\varphi_{0}\right]$, where $\varphi_{0}\equiv\hbar/2e$ is the reduced flux quantum.
@@
Since we are using node fluxes as our coordinates, we can write the “kinetic” energy in terms of the node fluxes’ time derivatives, which are equal to the classical node voltages. The energy contribution from a capacitor connecting nodes $j$ and $k$ is then $E=C_{jk}(\dot{\Phi}_{j}-\dot{\Phi}_{k})^{2}/2$, i.e., it depends on the difference of these node flux time derivatives across it. Combining the kinetic and potential energies we can write the Lagrangian $\mathcal{L}$. From this we can define the node charges as the canonical momenta, $q_{j}\equiv\partial\mathcal{L}/\partial\dot{\Phi_{j}}$, and write the classical Hamiltonian in terms of node charges and fluxes, plus bias voltages and fluxes. At this point we change charge/flux coordinates to operators, change to dimensionless operators $\hat{n}_{j}\equiv\hat{q}_{j}/2e$, $\hat{\phi_{j}}\equiv\hat{\Phi}_{j}/\varphi_{0}$ and obtain the quantum Hamiltonian as in Eq. 2. This procedure for defining node fluxes and charges ensures the canonical commutation relation $\left[\hat{\phi}_{j},\hat{n}_{j}\right]=i$.
@@
This process sounds daunting even when summarized. Luckily, the procedure is quite straightforward, if complicated, and has been both written about in great detail [41, 42] and has been integrated into open-source codes like scqubits, SQCircuit, and CircuitQ [44, 45, 46]. A user can input an arbitrary circuit, follow the procedure, and obtain a numerical solution of the eigenstates and energies. In practice, this requires truncating the charge and flux basis states so that the Hilbert space is finite and can be diagonalized numerically. However, for a circuit with any significant complexity, a complete numerical solution can be computationally infeasible even when bases are truncated. One solution is to perform hierarchical diagonalization [42, 47, 45]. In this approach, a designer uses their knowledge of a circuit to divide it into subcircuits, each of which has well-defined eigenstates that can be easily solved. The subcircuits are described using a basis of these eigenstates, which again is truncated at some finite dimension. The subcircuits are then coupled together and solved using the product space of these subcircuit basis states. This is similar to a perturbative approach but differs in that it performs exact diagonalization within subcircuits, allowing efficient treatment of degenerate or near-degenerate states without relying on small parameter expansions.
@@
In addition to the challenge of solving an arbitrary circuit, it is typically desirable to have an effective quantized Hamiltonian written in terms of oscillator creation/annihilation operators and Pauli matrices, as in Eq. 3. Such an effective Hamiltonian is more amenable to analytical theory and is usually more useful for quantum gate design and error correction architecture design. One solution is to use knowledge of the intended design to build up subcircuits that are weakly coupled—for instance, the coupled transmon and oscillator in Fig. 2 can be solved as an arbitrary circuit, but it is easier to recognize the transmon and oscillator modes and then add in the coupling in terms of these mode operators. Another approach is to exactly diagonalize the Hamiltonian, then write it in its eigenbasis with a set of Kerr, cross-Kerr, and higher-order terms as in Section II.3. This approach has the advantage of directly giving the energy spectrum, but is less useful if environmental conditions (flux or charge biases or drive tones) will be changed, as these can completely change the mode structure. Methods for rapidly calculating terms of interest such as inter-qubit couplings and dispersive couplings have been developed [48, 49], making the procedure less computational intensive.
@@
It is possible to connect nonlinear circuit elements in such a way that a Lagrangian cannot be defined via this procedure. For instance, a Josephson junction in parallel with a quantum phase slip junction has a Hamiltonian description, but deriving a consistent Lagrangian is nontrivial due to the nonlocal behavior. A phase slip junction is the dual of a Josephson junction, with Hamiltonian contribution $H_{Q}=-E_{Q}\cos\left(2\pi\frac{q}{2e}\right)$—the appearance of the charge in the nonlinear term breaks the standard procedure [50]. Likewise, the Lagrangian for a non-reciprocal device cannot be straightforwardly written down since the branch flux would depend on the direction we choose to calculate it [51]. Fortunately, frameworks for dealing with these situations have been formulated in geometric [52, 50, 53] and exactly quantized [54] terms as well as including open-systems effects [51]. We refer the reader to Ref. [53] for an excellent summary of the literature.
III.2 Open-systems effects
The Hamiltonian description of a circuit explained above does not include lossy effects. Here one may use a master equation or non-Hermitian Hamiltonian [55] to describe the circuit. However, in the most commonly-considered case, Purcell decay, a full quantum treatement is unnecessary. Purcell decay refers to the decay of a qubit via its coupling to a readout resonator, which itself is coupled to the lossy external microwave environment [5]. The term is sometimes also used to refer to decay to the external microwave environment due to any intended couplings, e.g., from qubit drive lines. In the case where the qubit and readout resonator are far detuned ($\Delta\equiv\omega_{R}-\omega_{Q}\gg g$) and the resonator has a high quality factor ($Q\equiv\omega_{R}/\kappa\gg 1$, where $\kappa$ is the resonator loss rate), the Purcell-induced decay rate is
$\Gamma_{1}^{\mathrm{Purcell}}=\frac{g^{2}}{\Delta^{2}}\kappa$ (13)
. This expression holds for a single resonator in the dispersive regime $g\ll\Delta$. When the resonator is driven with a coherent state with mean photon number $\bar{n}$ and the output is amplified with a phase-preserving amplifier, qubit state information is acquired at a rate
$\Gamma_{\mathrm{meas}}=\frac{\chi^{2}}{\chi^{2}+\kappa^{2}}\kappa\bar{n}$ (14)
[5]. We see here a tension: the measurement rate increases with $\kappa$ and $\chi$, but so does the Purcell decay rate (since $\chi\sim g^{2}$). The solution is to add a Purcell filter [56], which reduces the effective loss rate at the qubit frequency. Such a filter may be a bandpass near the resonator frequency or a bandstop near the qubit freuqency. This replaces $\kappa$ with $\kappa_{\mathrm{eff}}(\omega_{Q})$ in Eq. 13. Here $\kappa_{\mathrm{eff}}$ can be calculated as the linewidth of a resonator at $\omega_{R}$ that would produce the same real admittance as that seen by the qubit at $\omega_{Q}$. In the case of a bandpass filter at frequency $\omega_{F}\approx\omega_{R}$ with bandwidth $\kappa_{F}$, this yields
$\Gamma_{1}^{\mathrm{Purcell}}=\frac{g^{2}}{\Delta^{2}}\frac{\omega_{Q}}{\omega_{R}}\frac{\kappa_{F}}{2\Delta}\kappa$ (15)
where we have taken the limit that the filter bandwidth is much larger than the detuning between filter and resonator [57].
Note that this expression is exactly what one would obtain by treating the qubit as a classical harmonic oscillator and calculating $\Gamma_{1}=\mathrm{Re}\left\{Y(\omega_{Q})\right\}/C_{Q}$, where $Y$ is the admittance seen by the qubit and $C_{Q}$ is the total qubit capacitance. This classical admittance approach provides a convenient shortcut for transmon qubits which are only weakly anharmonic, and can function for any Purcell filter geometry. It also works for non-resonant couplings such as those to drive lines, and allows one to easily calculate the impact of the interference of qubit radiation with iteslf at different physical locations [58, 59]. For more anharmonic circuits circuits a prefactor $\beta\equiv\bra{0}\hat{n}\ket{1}/n_{01}^{QHO}$, i.e., the charge number matrix element at the node where the coupling occurs normalized by the harmonic oscillator charge matrix element, is required to account for the different effect of charge coupling on the circuit states. This is for the case where the qubit is capacitively coupled to the lossy mode or line; for inductive coupling we replace $\hat{n}\rightarrow\hat{\phi}$. Note that both these matrix elements are 1 for a harmonic oscillator circuit and near 1 for a transmon qubit.
Purcell decay is best addressed at the circuit graph level, as it is intrinsic to the couplings that are built into the graph and can be addressed as a microwave filtering problem. However, as we discuss in Section V.5.1, it can also be directly simulated from the device layout.
III.3 From Layout to Graph